### v1

# Testing stimulated emission photon direction

Abstract
While naively laser only allow to excite external target, for example STED microscopy demonstrates it can also cause its deexitation, however, under uncommon condition of being prepared as excited. These two causations are governed by stimulated emission-absorption pair of equations, and there are presented many arguments that they are switched in perspective of CPT symmetry, however, it would mean photon direction opposite to usually assumed. This article discusses these arguments and proposes simple STED-like tests to experimentally determine this direction once for all, to be performed in a near future. Depending on its result: it should lead to new applications like 2WQC more symmetric quantum computers, or macroscopic violation of CPT symmetry, so far tested only in microscopic settings. Also forward/backward/bidirectional photon sources are discussed.
Keywords: lasers, photonics, STED microscopy, CPT symmetry, ring laser, synchrotron radiation, 2WQC
I Introduction
CPT symmetry (of charge conjugation (C) + parity transformation (P) + time reversal (T)) is shown [1] to be necessary for local Lorentz invariant quantum field theories, hence is generally believed to be satisfied by our nature. It has dozens of successful experimental confirmations [2], however, they are microscopic - come from observation of single particles. Seems conscious macroscopic tests were not performed so far, interesting especially due to T time symmetry which might lead to new applications - some are proposed and discussed here.
As in Fig. 1, applying CPT symmetry to scenario ”laser causes target excitation”, it becomes ”CPT(laser) causes CPT(target) deexcitation”, replacing absorption equation with stimulated emission, with photons traveling from CPT(target) to CPT(laser). In contrast, standard textbook view on stimulated emission is energy being kind of knocked by incoming photon - with the opposite direction for the emitted photon.
As there are at least two possibilities for photon direction, the final judge should be experiment. Related tests were earlier proposed using free electron laser [3], or ring laser [4]. However, they turned out too difficult to organize. This article simplifies test setting to STED-like with just two e.g. diode lasers, determining photon direction through delay. Hence it should be organized in a near future - to update this article and depending on the outcome: leading to further mentioned novel applications, or suggesting macroscopic violation of CPT symmetry.
There are also gathered novel arguments, and proposed view on photon sources: using forward/backward/bidirectional beam as in Fig. 2: applying to external target absorption/stimulated emission for unidirectional, or both equations for bidirectional.
II STED microscopy based tests
Stimulated emission depletion (STED) microscopy [5] improves resolution by reducing size of excited dot with donought-like depletion beam from second laser as in Fig. 2. It uses fluorescent dye being approximately 3 level medium: excitation with one laser from 0 to 2, then fast relaxation to 1, then deexcitation from 1 to 0 with slightly lower wavelength - usually through just spontaneous emission observed by microscope lens.
In STED we additionally speedup this deexcitation through stimulated emission with additional laser of frequency for transition - reducing the number of photons from this region observed by microscope lens focused on spontaneous emission, this way allowing to increase microscope resolution.
While spontaneous emission photons should have a random direction, the question of concern is direction of such stimulated emission photons: toward or outward the laser? Textbooks assume the latter, discussion here based on CPT symmetry suggests the former. There might be also a different angle, what also could be tested experimentally.
II-A Delay STED-like test
Assuming there are only these two options: or , the simplest way to distinguish them seems delay. There is literature discussing delay in STED microscopy, like [6, 7, 8], but they seem insufficient to uniquely answer the direction question.
Therefore, there is proposed STED-like setting as in the bottom of Fig. 1: two (e.g. diode) lasers: continuous for excitation of the dye, and impulse for its stimulated emission. There is also photodetector observing spontaneous emission from this dye, which should observe reduced intensity due to depletion laser impulse - the question is its delay.
Hence photodetector should be connected to oscilloscope, triggered by depletion laser impulse. Its observed delay contains positive contribution for light path from dye to detection, and contribution from laser-dye path, of sign for one of the two assumed possibilities - which could be determined from this delay, e.g. from oscilloscope snapshots planned to be included in the next version of this article.
II-B Alternative STED-like tests
Another approach is placing photodetector in depletion laser-dye line behind the dye (yellow if Fig. 1), allowing to determine or from intensities. Measuring also angle dependence, we could also test (rather exclude) change of angle hypothesis.
As discussed further, additional tests could be using unidirectional e.g. ring laser: its forward beam should not allow for depletion here, its backward beam should allow - what could be also tested in the future.
III CPT symmetry in photonics arguments
Let us gather here some arguments for the expected possibility mainly based on CPT symmetry. Test pointing instead, would need to overcome these arguments, e.g. through macroscopic CPT violation.
III-A Bidirectional and forward/backward beams
Stimulated emission-absorption equations are at heart of lasers, governing deexcitation-excitation e.g. in laser active medium. For atoms in the ground state, excited:
where are (CPT symmetric) Einstein’s coefficients [11], is radiation density of the incident field at frequency , corresponding to transition between the two considered states: .
While both are assumed to act on active medium inside laser, naively only absorption equation acts on external target - for example STED laser demonstrates it is not true, acting e.g. by diode laser with stimulated emission equation on external target. However, this equation is negligible for - requires initial excitation, while unprepared target is close to the ground state.
As in Fig. 2, the situation changes for example for ring laser with optical isolator - using Faraday effect [12] to allow photon circulation only in one direction. Such photons could leave the laser and be absorbed by a (forward) external target.
Looking at such unidirectional ring laser from CPT perspective, photon trajectory would be reversed, causing excitation of different (backward) target - increasing toward minus time, hence decreasing in standard time perspective - as in stimulated emission.
Therefore, this CPT symmetry switches the two equations - in contrast e.g. to diode lasers in STED microscopy, seems only one equation at the time acts on the forward/backward targets - allowing to split them into intuitively forward/backward beam, combined into bidirectional beam in standard e.g. diode laser.
Another example of laser sources switching targets in CPT perspective are based on synchrotron radiation like free electron laser/synchrotron/wiggler/undulator - while there is usually focus only on their forward target, in CPT perspective electron trajectories would be reversed - causing excitation of backward target in the opposite direction, hence should reduce its in standard time perspective, as in stimulated emission.
While application of CPT symmetry seems abstract, general relativity in theory allows a more tangible possibility useful for intuitions - in theory allowing Klein-bottle-like nonorientable wormholes [10] applying P or T symmetry to e.g. a rocket travelling through which - for T symmetry switching absorption/stimulated emission inside, forward/backward beams.
III-B Experimental consequences like Rabi cycle
To test the above with STED-like setting, as mentioned we can replace depletion laser with forward/backward beam - only backward should lead to depletion. If stimulated emission-absorption use opposite photon trajectories, we could also try to split them exploiting delay with fast shutter or switch.
Another type of consequence is population in external target of laser. For bidirectional acting with both equations, we could at most reach population, and it is indeed well known as the final population of damped Rabi cycle [13]: using periodically stimulated emission and absorption equations. In Fig. 3 there is suggested hydrodynamic analog of Rabi cycle (another can be found in [14]), which should exist through EM-hydrodynamics analogy as written, like the shown resonator coupled with external object exchanging momentum in both directions, what might be also appropriate intuition for bidirectional lasers, using cyclically positive and negative radiation pressure [16]. Rabi cycle in CPT perspective should be practically the same - bringing another argument for photons travelling there only between laser and target.
However, using only forward beam, we should be able to reach nearly complete excitation, and such effect has turned out to be known in literature: despite many attempts and theoretical prediction, they were not able to obtain Rabi oscillation with free electron laser [9] - instead getting Autler-Townes effect (dynamics Stark) leading to , instead of for damped Rabi oscillations.
III-C Some possible new applications
If confirmed, the above considerations could lead to many new applications, for example of currently neglected backward beam, like:
Forward beam could allow for state preparation of quantum computers to , backward beam for its CPT analog (postparation) for 2WQC: two-way quantum computers [4].
Backward beam for attacks on quantum cryptographic protocols - use external laser to additionally stimulate emission inside communication laser to steal/intercept and measure such additional photons.
Backward beam for 3D printing/reversed lithography: excite material and precisely stimulate its deexcitation - controlling position and time, also frequency for multiple possible transitions.
Synchrotron sources usually have much higher photon energy, its backward beam could increase probability of some e.g. particle events, which CPT transform use such photons coming from e.g. synchrotron.
IV Conclusions and further work
There were proposed various tests for photon direction of stimulated emission, one of them should be soon performed to update this article. There were also presented many arguments based on CPT symmetry that direction should be from target to laser, which is the opposite than textbook explanation.
The further work depends on the result of the proposed test:
as required by CPT symmetry should lead to many new applications to explore, like the one listed above,
would indicate macroscopic violation of CPT symmetry, requiring further investigation of its details.

### v3

@@
 Abstract
-While naively laser only allow to excite external target, for example STED microscopy demonstrates it can also cause its deexitation, however, under uncommon condition of being prepared as excited.
-These two causations are governed by stimulated emission-absorption pair of equations, and there are presented many arguments that they are switched in perspective of CPT symmetry, however, it would mean photon direction opposite to usually assumed.
-This article discusses these arguments and proposes simple STED-like tests to experimentally determine this direction once for all, to be performed in a near future.
-Depending on its result: it should lead to new applications like 2WQC more symmetric quantum computers, or macroscopic violation of CPT symmetry, so far tested only in microscopic settings.
-Also forward/backward/bidirectional photon sources are discussed.
+While naively laser only causes excitation of external target, for example STED microscopy demonstrates it can also cause its deexitation, however, under uncommon condition of being prepared as excited.
+These two causalities are governed by stimulated emission-absorption pair of equations, and swap places in perspective of CPT symmetry, however, it means photon direction of stimulated emission should be opposite to usually assumed.
+This article discusses various arguments and proposes simple STED-like tests to experimentally determine stimulated emission photon direction(s) once for all, possibly both, to be performed in a near future, already having indirect experimental evidence, like lack of Rabi cycle for free electron laser, or recent ”negative time evidence” experiment in the news.
+Depending on its result: it should lead to new applications like 2WQC more symmetric quantum computers, or macroscopic violation of CPT symmetry, so far probably tested only in microscopic settings.
+Also proposed forward/backward/bidirectional photon sources are discussed.
 Keywords: lasers, photonics, STED microscopy, CPT symmetry, ring laser, synchrotron radiation, 2WQC
@@
 CPT symmetry (of charge conjugation (C) + parity transformation (P) + time reversal (T)) is shown [1] to be necessary for local Lorentz invariant quantum field theories, hence is generally believed to be satisfied by our nature.
-It has dozens of successful experimental confirmations [2], however, they are microscopic - come from observation of single particles.
-Seems conscious macroscopic tests were not performed so far, interesting especially due to T time symmetry which might lead to new applications - some are proposed and discussed here.
+It has dozens of successful experimental confirmations [2], however, it seems all tests so far were microscopic - come from observations of single particles.
+Conscious macroscopic test was not found in literature, important especially due to T time symmetry which might lead to new applications - some are proposed and discussed here.
+For example through system response before impulse, like recently experimentally observed [3] in the news, for example in Scientific American 111 www.scientificamerican.com/article/evidence-of-negative-time-found-….
 As in Fig.
-1, applying CPT symmetry to scenario ”laser causes target excitation”, it becomes ”CPT(laser) causes CPT(target) deexcitation”, replacing absorption equation with stimulated emission, with photons traveling from CPT(target) to CPT(laser).
+1, 2, applying CPT symmetry to scenario ”laser causes target excitation”, it becomes ”CPT(laser) causes CPT(target) deexcitation”, replacing absorption equation with stimulated emission, with photons traveling from CPT(target) to CPT(laser).
 In contrast, standard textbook view on stimulated emission is energy being kind of knocked by incoming photon - with the opposite direction for the emitted photon.
+As in Fig.
+2, the above consequence of CPT symmetry also allows to conclude the textbook view, hence both can be true.
 As there are at least two possibilities for photon direction, the final judge should be experiment.
-Related tests were earlier proposed using free electron laser [3], or ring laser [4].
+Related tests were earlier proposed using free electron laser [4] (earlier 2009 post), or ring laser [5].
 However, they turned out too difficult to organize.
 This article simplifies test setting to STED-like with just two e.g.
-diode lasers, determining photon direction through delay.
-Hence it should be organized in a near future - to update this article and depending on the outcome: leading to further mentioned novel applications, or suggesting macroscopic violation of CPT symmetry.
+diode lasers, fluorescent dye and fast photodetector - determining photon direction through delay.
+Hence it should be organized soon - to update this article and depending on the outcome: leading to further discussed novel applications, or indicating macroscopic violation of CPT symmetry.
 There are also gathered novel arguments, and proposed view on photon sources: using forward/backward/bidirectional beam as in Fig.
-2: applying to external target absorption/stimulated emission for unidirectional, or both equations for bidirectional.
+3: applying to external target absorption/stimulated emission for unidirectional, or both equations for bidirectional.
+Updated article version was motivated by ”negative time evidence” in the news [3], which observed clear signal both before and after the impulse, already indicating absorption-stimulated emission causalities of opposite time delays, already bringing some CPT symmetry suggestions.
+Also, QInter of our 2WQC group (https://www.qaif.org/2wqc): Vysakh Saji has recently performed the proposed experiment, obtaining unofficial result as ns for setting, this way confirming validity of CPT symmetry in macroscopic scale, to be updated with official results.
+Second update observes that textbook view can be also concluded from CPT symmetry hence both photon directions can be true, also discusses transparency.
 II STED microscopy based tests
 Stimulated emission depletion (STED) microscopy [6] improves resolution by reducing size of excited dot with donought-like depletion beam from second laser as in Fig.
-2.
-It uses fluorescent dye being approximately 3 level medium: excitation with one laser from 0 to 2, then fast relaxation to 1, then deexcitation from 1 to 0 with slightly lower wavelength - usually through just spontaneous emission observed by microscope lens.
-In STED we additionally speedup this deexcitation through stimulated emission with additional laser of frequency for transition - reducing the number of photons from this region observed by microscope lens focused on spontaneous emission, this way allowing to increase microscope resolution.
+3.
+It uses fluorescent dye being approximately 3 level medium: excited with one laser from 0 to 2, then fast relaxation to 1, then deexcitation from 1 to 0 with slightly lower wavelength - usually through just spontaneous emission observed by microscope lens.
+In STED we additionally speedup this deexcitation through stimulated emission by second: depletion laser of frequency for transition - reducing the number of photons from this region monitored by microscope lens focused on spontaneous emission, this way allowing to increase microscope resolution.
 While spontaneous emission photons should have a random direction, the question of concern is direction of such stimulated emission photons: toward or outward the laser?
 Textbooks assume the latter, discussion here based on CPT symmetry suggests the former.
-There might be also a different angle, what also could be tested experimentally.
+There can be also both, or a different angle, what also could be tested experimentally.
 II-A Delay STED-like test
 Assuming there are only these two options: or , the simplest way to distinguish them seems delay.
-There is literature discussing delay in STED microscopy, like [6, 7, 8], but they seem insufficient to uniquely answer the direction question.
+There is literature discussing delay in STED microscopy, like [7, 8, 9], however, data they provide seem insufficient to uniquely answer this direction question.
 Therefore, there is proposed STED-like setting as in the bottom of Fig.
 1: two (e.g.
-diode) lasers: continuous for excitation of the dye, and impulse for its stimulated emission.
+diode) lasers: continuous for excitation of the dye, and impulse depletion for its stimulated emission.
 There is also photodetector observing spontaneous emission from this dye, which should observe reduced intensity due to depletion laser impulse - the question is its delay.
@@
 from oscilloscope snapshots planned to be included in the next version of this article.
+Fluorescent dye can be replaced with a different medium for which we can maintain a relatively high population level of excited atoms, what is much more accessible for 3,4-level media.
+Impulse preferably should have a fraction of nanosecond, (e.g.
+cascade) photodetector needs to allow nanosecond-scale time resolution.
+However, their imperfections could be compensated by increasing distance to impulse laser.
 II-B Alternative STED-like tests
-Another approach is placing photodetector in depletion laser-dye line behind the dye (yellow if Fig.
-1), allowing to determine or from intensities.
+Another approach is placing photodetector in depletion laser-dye line behind the dye, allowing to determine or from intensities.
 Measuring also angle dependence, we could also test (rather exclude) change of angle hypothesis.
@@
 ring laser: its forward beam should not allow for depletion here, its backward beam should allow - what could be also tested in the future.
+We could also test suggestions of better transparency for backward beam, e.g.
+by putting (non-excited) obstacle on the way to depletion laser in STED-like setting.
+It also brings hope for effects on excited targets already for shielded synchrotron sources.
+Even if such photons could not cross the obstacle, this effect could be used e.g.
+for information transfer by trying to detect such additional directional photons from excited target behind obstacle, or monitoring its population level.
+Rabi cycle could be used instead, by applying impulse laser and monitoring/analyzing situation at impulse boundaries.
 III CPT symmetry in photonics arguments
-Let us gather here some arguments for the expected possibility mainly based on CPT symmetry. Test pointing instead, would need to overcome these arguments, e.g. through macroscopic CPT violation.
+Let us gather here some arguments for the expected possibility mainly based on CPT symmetry.
+However, as in Fig.
+2, it seems to also imply textbook case, hence both can be true - practically this is test of existence of effect.
 III-A Bidirectional and forward/backward beams
@@
 As in Fig.
-2, the situation changes for example for ring laser with optical isolator - using Faraday effect [12] to allow photon circulation only in one direction.
+3, the situation changes for example for ring laser with optical isolator - using Faraday effect [11] to allow photon circulation only in one direction.
 Such photons could leave the laser and be absorbed by a (forward) external target.
@@
 Therefore, this CPT symmetry switches the two equations - in contrast e.g.
-to diode lasers in STED microscopy, seems only one equation at the time acts on the forward/backward targets - allowing to split them into intuitively forward/backward beam, combined into bidirectional beam in standard e.g.
+to diode lasers in STED microscopy, seems only one equation at the time acts on the forward/backward targets - allowing to split beams them into intuitively forward/backward beam acting with separated absorption/stimulated emission equation alone, combined into bidirectional beam in standard e.g.
 diode laser.
 Another example of laser sources switching targets in CPT perspective are based on synchrotron radiation like free electron laser/synchrotron/wiggler/undulator - while there is usually focus only on their forward target, in CPT perspective electron trajectories would be reversed - causing excitation of backward target in the opposite direction, hence should reduce its in standard time perspective, as in stimulated emission.
+As noticed in Fig.
+2, forward/backward beams might have completely different transparency as they act only on deexcited / excited atoms.
+As everyday matter usually has , backward beam might have much better penetration, allowing for novel applications.
 While application of CPT symmetry seems abstract, general relativity in theory allows a more tangible possibility useful for intuitions - in theory allowing Klein-bottle-like nonorientable wormholes [12] applying P or T symmetry to e.g.
@@
 To test the above with STED-like setting, as mentioned we can replace depletion laser with forward/backward beam - only backward should lead to depletion.
-If stimulated emission-absorption use opposite photon trajectories, we could also try to split them exploiting delay with fast shutter or switch.
+If stimulated emission-absorption use opposite photon trajectories, we could also try to split them exploiting delay with fast shutter or switch, or maybe exploiting transparency absorbing only forward beam.
 Another type of consequence is population in external target of laser.
@@
 III-C Some possible new applications
-If confirmed, the above considerations could lead to many new applications, for example of currently neglected backward beam, like:
-Forward beam could allow for state preparation of quantum computers to , backward beam for its CPT analog (postparation) for 2WQC: two-way quantum computers [4].
-Backward beam for attacks on quantum cryptographic protocols - use external laser to additionally stimulate emission inside communication laser to steal/intercept and measure such additional photons.
-Backward beam for 3D printing/reversed lithography: excite material and precisely stimulate its deexcitation - controlling position and time, also frequency for multiple possible transitions.
-Synchrotron sources usually have much higher photon energy, its backward beam could increase probability of some e.g.
-particle events, which CPT transform use such photons coming from e.g.
+If confirmed, the above considerations could lead to many new applications, especially of currently neglected backward beam, like:
+If we could send information this way, there could be e.g.
+built time-loop computers (Section V of [4]) for example using backward beam to send chip output back to its original input with negative time delay, also in theory allowing to solve NP problems.
+Connecting multiple such negative time transmitters e.g.
+into a cycle, could allow to also reach macroscopic negative time differences.
+Backward camera e.g.
+exploiting transparency or for astronomical observations - focused on target’s backward beam, while standard camera is focused on its forward beam.
+It would replace CMOS/CCD matrix behind the lens, with matrix of continuously excited e.g.
+fluorescent dye, plus photodetectors monitoring their spontaneous emission.
+Backward beam should be transparent for medium, what might allow for much better penetration than forward (transparent for ) - for example for backward camera or communication through obstacles.
+Backward beam for attacks on quantum cryptographic protocols like BB84 [18] - use external laser to additionally stimulate emission inside communication laser to steal/intercept such additional photons, which could be read e.g.
+by passing through control of CNOT gate - affecting second qubit to be measured.
+Backward beam for observations actively increasing deexcitation rate of excited objects, e.g. for thermal imaging,
+Backward beam for 3D printing/reversed lithography: excite material and precisely stimulate its deexcitation - controlling position and time, also frequency to choose one of multiple possible transitions, e.g.
+some high energy transition to locally reduce temperature.
+Synchrotron sources usually have much higher photon energy, its backward beam could e.g.
+increase probability of selected nuclear decays/transitions: producing photons of chosen energy.
+Maybe also for other particle events: which CPT transform use such photons coming from e.g.
 synchrotron.
 IV Conclusions and further work
-There were proposed various tests for photon direction of stimulated emission, one of them should be soon performed to update this article.
-There were also presented many arguments based on CPT symmetry that direction should be from target to laser, which is the opposite than textbook explanation.
+There were proposed various tests for photon direction of stimulated emission, soon to be performed to update this article.
+There were also presented many arguments based on CPT symmetry that direction should be from target to laser, which is the opposite than textbook explanation, but allowing to imply it as a consequence.
+Analogous processes could be also possible for other EM frequencies like radio to be explored.
 The further work depends on the result of the proposed test:
 as required by CPT symmetry should lead to many new applications to explore, like the one listed above,
-would indicate macroscopic violation of CPT symmetry, requiring further investigation of its details.
+lack of would indicate macroscopic violation of CPT symmetry, requiring further investigations of its details.